OpenStudy (anonymous):

solve the following quadratic equation x2+2x-15=0

OpenStudy (anonymous):

x^2 + 5x - 3x - 15 = 0 x(x+5) - 3(x+5) = 0 (x+5) (x-3) = 0 x = -5, 3

OpenStudy (anonymous):

i s0lve this question same like that but teacher said its wrong ans :(

OpenStudy (anonymous):

But this is the right answer.

OpenStudy (anonymous):

hmmm thn why teacher said its wrong ans :/

OpenStudy (anonymous):

The you need to ask the teacher for the solution. But this is right

OpenStudy (anonymous):

can u help me in understand t0 matrix?

OpenStudy (anonymous):

what help you need?

OpenStudy (anonymous):

how we find inverse matrix?

OpenStudy (anonymous):

reply meeeeeee????

OpenStudy (anonymous):

Read this.. http://mathworld.wolfram.com/MatrixInverse.html this article will give you a good insight about inverse matrix